# AortaDiff: A Unified Multitask Diffusion Framework for Contrast-Free AAA Imaging

### v1

@@ 4.5 Baselines
Synthetic CECT Generation. To evaluate the performance of our method on synthetic CECT generation, we consider two baseline architectures: a U-Net [26] and a conditional diffusion model (CDM) [27]. The noise prediction network used in all diffusion models, including our proposed method, shares the same U-Net backbone for a fair architectural comparison. For the CDM baseline, we evaluate two variants:
@@
This comparison allows us to assess whether focusing generation on the anatomically relevant region yields better results than generating the entire image.
Lumen Segmentation. For the lumen segmentation task, we compare our method with nnU-Net [28], a state-of-the-art self-configuring medical image segmentation framework that achieves strong performance across a wide range of datasets. Additionally, we include MedSAM [29], a pretrained vision foundation model for medical image segmentation. For MedSAM, we use the publicly released pretrained weights without any task-specific fine-tuning.
@@ 4.6.1 Quantitative Evaluation of CECT Synthesis
Table 1 presents quantitative results for CECT synthesis task, evaluated within the local aortic region. We observe that our proposed aorta-restricted generation strategy (CDM-A) outperforms both CDM-W and U-Net across all metrics. While diffusion models generally yield higher-fidelity outputs than U-Net, CDM-W, which conditions on the entire NCCT slice, underperforms U-Net. This performance gap highlights the diffusion model’s sensitivity to irrelevant regions in the conditioning input, such as those introduced by imperfect alignment between NCCT and CECT scans. These findings underscore the necessity of region-of-interest-focused generation to ensure stable performance in anatomically localized tasks.
@@ 4.6.2 Quantitative Evaluation of Lumen and Thrombus Segmentation
The results reveal several key findings. First, for generation-only method, CDM-A outperforms both CDM-W and U-Net in both tasks, confirming that ROI-constrained generation improves downstream segmentation quality. Second, AortaDiff-P achieves the highest Dice scores across both tasks, 0.8933 for lumen and 0.5326 for thrombus, exceeding even fully supervised segmentation-only methods.
@@ 4.7 Clinical Measurements for AAA
Aneurysm morphology is important for predicting AAA progression [13, 30]. We evaluate two clinically relevant metrics, lumen diameter and thrombus area, on a random sample of ten test cases. As shown in Tables 3 and 4, our models produce accurate predictions on both metrics and outperform nnU-Net, indicating strong potential for clinical application.
@@ 5 Conclusions
In this work, we introduced AortaDiff, a unified multi-task diffusion framework for generating synthetic CECTs and segmenting the aortic lumen directly from NCCT scans. We demonstrated that having a diffusion model perform these tasks simultaneously enhances performance across both synthesis and segmentation. We also showed that an aorta-focused generation strategy with a background fusion technique can overcome significant data misalignment issues, while a direct, semi-supervised training scheme enables effective learning from annotation-scarce clinical datasets. Experimentally, our multitasking approach significantly outperformed state-of-the-art single-task and multi-stage baselines on both synthesis and segmentation, leading to more accurate AAA clinical measurements. Ultimately, this research represents a meaningful step towards creating a “digital contrast” for CT. By reducing the reliance on iodinated contrast agents, such AI-driven workflows can enhance patient safety, lower procedural costs, and minimize the environmental impact of medical waste, contributing to a more sustainable and climate-neutral healthcare system.

### v2

@@ 3.5 Baselines
Synthetic CECT Generation. To evaluate the performance of our method on synthetic CECT generation, we consider three baseline architectures: a U-Net [unet], a CycleGAN[zhu2020unpairedimagetoimagetranslationusing] and a conditional diffusion model (CDM) [adm]. The noise prediction network used in all diffusion models, including our proposed method, shares the same U-Net backbone for a fair architectural comparison. The CycleGAN experiments follow the default configurations provided in the official codebase. For the CDM baseline, we evaluate two variants:
@@
This comparison allows us to assess whether focusing generation on the anatomically relevant region yields better results than generating the entire image. For the U-Net and CycleGAN baselines, the conditioning input is the whole NCCT image.
Due to misalignment between NCCT and CECT images, we replace the default MSE loss with a perceptual loss for all baseline models. In our ablation study, we demonstrate that incorporating perceptual loss improves performance across all models.
Lumen Segmentation. For the lumen segmentation task, we compare our method with nnU-Net [nnunet], a state-of-the-art medical image segmentation framework that achieves strong performance. Additionally, we include MedSAM [medsam], a pretrained vision foundation model for medical image segmentation. For MedSAM, we use the publicly released pretrained weights without any task-specific fine-tuning.
@@ 3.6.1 Quantitative Evaluation of CECT Synthesis
Table 1 presents quantitative results for CECT synthesis task, evaluated within the local aortic region. We observe that our proposed aorta-restricted generation strategy (CDM-A) outperforms both CDM-W, CycleGAN and U-Net across all metrics. CDM-W and CycleGAN underperform U-Net. This performance gap highlights the sensitivity of diffusion and GAN models to irrelevant regions in the conditioning input, such as those introduced by imperfect alignment between NCCT and CECT scans. These findings underscore the necessity of region-of-interest-focused generation to ensure stable performance in anatomically localized tasks.
@@ 3.6.2 Quantitative Evaluation of Lumen and Thrombus Segmentation
The results reveal several key findings. First, for generation-only method, CDM-A outperforms both CDM-W, CycleGAN and U-Net in both tasks, confirming that ROI-constrained generation improves downstream segmentation quality. Second, AortaDiff-P achieves the highest Dice scores across both tasks, 0.8933 for lumen and 0.5326 for thrombus, exceeding even fully supervised segmentation-only methods.
@@ 3.7 Clinical Measurements for AAA
Aneurysm morphology is important for predicting AAA progression [oxaaa, Wanhainen2018-nk]. We evaluate two clinically relevant metrics, lumen diameter and thrombus area, on a random sample of ten test cases. As shown in Tables 3 and 4, our models produce accurate predictions on both metrics and outperform nnU-Net, indicating strong potential for clinical application.
@@
4 Ablation Study
We conducted an ablation study to justify our design choices and provide insight into how these designs influence the results.
We evaluate the following design choices:
  • Reconstruction Loss: use MSE or perceptual loss for CECT image reconstruction.
  • Mask representation: use signed distance function to represent lumen mask instead of binary mask.
  • Multi-task loss combination: use the uncertainty-aware multi-task loss instead of using the equal-weighted sum of segmentation and generation loss.
From the reconstruction results in Table 5, we observe that using LPIPS loss yields better reconstruction performance compared to MSE loss across all three models. This improvement is likely due to the background misalignments between our NCCT–CECT pairs, where a pixel-wise loss forces the model to learn these misalignments instead of focusing on perceptually meaningful features. For both multi-task models, AortaDiff-F and AortaDiff-P, the results in Table 6 show that using an SDF-based mask representation yields better segmentation performance compared to a binary mask. In addition, Table 7 shows that using the uncertainty-aware loss leads to improved performance for both tasks compared to the equal-weight loss formulation.
Table 5: Ablation results comparing different reconstruction losses, using either MSE or perceptual loss.
Reconstruction Loss PSNR (dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CDM with MSE loss 20.87 0.7253 0.1077 CDM with LPIPS loss 23.80 0.7616 0.0859 CycleGAN with MSE loss 22.83 0.6492 0.1283 CycleGAN with LPIPS loss 23.19 0.6535 0.1129 U-Net with MSE loss 23.83 0.7602 0.0971 U-Net with LPIPS loss 24.40 0.7785 0.0815
Table 6: Ablation results for different loss combination strategies.
Mask Representation PSNR (dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Lumen Dice $\uparrow$ AortaDiff-F Binary 25.41 0.8262 0.0688 0.8753 AortaDiff-F SDF 25.61 0.8385 0.0671 0.8887 AortaDiff-P Binary 25.47 0.8198 0.0661 0.8760 AortaDiff-P SDF 25.48 0.8296 0.0606 0.8933
Table 7: Ablation results for multi-task loss. All metrics are the same as in the main paper. PSNR, SSIM, and LPIPS are computed over the aorta region.
Multi-task Loss PSNR (dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Lumen Dice $\uparrow$ AortaDiff-F with Equal-Weight Loss 25.37 0.8207 0.0660 0.8700 AortaDiff-F with Uncertainty-Aware Loss 25.61 0.8385 0.0671 0.8887 AortaDiff-P with Equal-Weight Loss 24.40 0.7808 0.0757 0.8830 AortaDiff-P with Uncertainty-Aware Loss 25.48 0.8296 0.0606 0.8933
@@ 5 Conclusions
In this work, we introduced AortaDiff, a unified multi-task diffusion framework for generating synthetic CECTs and segmenting the aortic lumen directly from NCCT scans. We demonstrated that having a diffusion model perform these tasks simultaneously enhances performance across both synthesis and segmentation. We also showed that an aorta-focused generation strategy with a background fusion technique can overcome significant data misalignment issues, while a direct, semi-supervised training scheme enables learning from annotation-scarce clinical datasets. Experiments show that our multitasking approach outperformed state-of-the-art single-task and multi-stage baselines on both synthesis and segmentation, leading to more accurate AAA clinical measurements. This research represents a meaningful step towards creating a “digital contrast” for CT.
Despite these contributions, limitations remain. First, DDPM-based models have longer inference time than CycleGAN or U-Net based models, and the use of RePaint further increases inference time, posing challenges for clinical deployment. Second, our method currently requires lumen masks during training; since the model learns to generate the lumen, future research could investigate leveraging latent representations for unsupervised lumen prediction. Third, uncertainty estimation is crucial for clinical trust, motivating future work on quantifying uncertainty in both generation and segmentation.